# Auto-Compressing Networks

### v1

@@ Abstract
Deep neural networks with short residual connections have demonstrated remarkable success across domains, but increasing depth often introduces computational redundancy without corresponding improvements in representation quality. In this work, we introduce Auto-Compressing Networks (ACNs), an architectural variant where additive long feedforward connections from each layer to the output replace traditional short residual connections. By analyzing the distinct dynamics induced by this modification, we reveal a unique property we coin as auto-compression—the ability of a network to organically compress information during training with gradient descent, through architectural design alone. Through auto-compression, information is dynamically "pushed" into early layers during training, enhancing their representational quality and revealing potential redundancy in deeper ones, resulting in a sparse yet powerful network at inference. We theoretically show that this property emerges from layer-wise training patterns present in ACNs, where layers are dynamically utilized during training based on task requirements. We also find that ACNs exhibit enhanced noise robustness compared to residual networks, superior performance in low-data settings, improved transfer learning capabilities, and mitigate catastrophic forgetting suggesting that they learn representations that generalize better despite using fewer parameters. Our results demonstrate up to 18% reduction in catastrophic forgetting and 30-80% architectural compression while maintaining accuracy across vision transformers, MLP-mixers, and BERT architectures. Furthermore, we demonstrate that when coupling ACNs with traditional pruning techniques, the compression gain persists and enables significantly better sparsity-performance trade-offs compared to conventional architectures. These findings establish ACNs as a practical approach to developing efficient neural architectures that automatically adapt their computational footprint to task complexity, while learning robust representations suitable for noisy real-world tasks and continual learning scenarios.
@@ 1 Introduction
Table 1: Connectivity, Forward and Backward Propagation for FFN, ResNet, and ACN architectures.
Deep learning has achieved significant breakthroughs across diverse tasks and domains (Krizhevsky et al., 2012; LeCun et al., 2015; Brown et al., 2020); however, it still lacks the flexibility, robustness, and efficiency of biological networks. Modern models rely on deep architectures with billions of parameters, leading to high computational, storage, and energy costs. Architecturally, these large models are primarily characterized by short residual connections (He et al., 2016), a design initially developed to enable robust training of deep neural networks via backpropagation. These skip connections establish a network topology where multiple information pathways are created (Veit et al., 2016), resulting in an ensemble-like behavior that delivers more efficient training and superior generalization compared to traditional feedforward networks.
Historically, since the emergence of Highway Networks (Srivastava et al., 2015), which first proposed additive skip connections researchers have explored numerous architectural variations. Residual Networks (ResNets) (He et al., 2016) removed learned gating functions and adopted direct identity skip connections becoming the industry standard. DenseNets (Huang et al., 2017b) utilized feature concatenation instead of addition, while FractalNets (Larsson et al., 2016) introduced a recursive tree-like architecture combining subnetworks of multiple depths to further enrich feature fusion. More recent works include learned weighted averaging across layer outputs (Pagliardini et al., 2024), application of attention mechanisms across block outputs (ElNokrashy et al., 2022) and denser connectivity patterns between network nodes (Zhu et al., 2025). Other works have explored adding scalars to either the residual or block stream to improve performance, training stability and representation learning (Savarese et al., 2016; Bachlechner et al., 2021; Zhang et al., 2024; Fischer et al., 2023). In neural machine translation, researchers have drawn inspiration from both vision and language domains to combine information from different layers, enabling richer semantic and spatial propagation throughout the network (Dou et al., 2018; Yu et al., 2018).
@@
While Residual Networks, as discussed, offer numerous advantages—such as more robust training and improved generalization—there remain several aspects that are worth some further examination and discussion. As highlighted in (Veit et al., 2016), these architectures exhibit a notable resilience to layer dropping and permutation. In (Huang et al., 2016), it was further observed that dropping subsets of layers during training can reduce overfitting and improve generalization. In a related study, (Alain & Bengio, 2016) showed that introducing skip connections between layers can lead to parts of the network being effectively bypassed and under-trained. More recently, research has revealed substantial parameter redundancy in large-scale foundation models, particularly within their deeper layers (e.g., (Gromov et al., 2024)). All these observations can be unified under the perspective that, although residual architectures facilitate training via multiple signal pathways, these same pathways can sometimes act as shortcuts that cause certain components to be either underutilized or prone to overfitting—ultimately limiting effective generalization. Supporting this concern, (Zhang et al., 2024) demonstrated that unscaled residual connections can degrade the quality of generative representation learning, offering a concrete case where standard (unscaled) residual connections negatively impact performance. Thus, an open question remains: can we design in a principled way alternative architectures that retain the key benefits of Residual Networks—such as multiple signal pathways and efficient gradient flow—while mitigating drawbacks such as potential redundancy and shortcut overuse, effectively resulting in better representation learning?
@@
Despite the breadth of the discussed research exploring different connectivity patterns across domains—with goals ranging from improved expressivity and representation learning to increased training efficiency—none of these potential improvements have achieved broad adoption beyond standard residual connections. In this work, moving a step towards answering the above questions, we explore an architectural variant where additive long feedforward connections from each layer to the output replace traditional short residual connections as shown in Table 1, introducing Auto-Compressing Networks (ACNs). ACNs showcase a unique property we coin as auto-compression—the ability of a network to organically compress information during training with gradient descent, through architectural design alone, dynamically pushing information to bottom layers, enhancing their representational quality, and naturally revealing redundant in deeper layers. We theoretically investigate the emergence of this property by analyzing the gradient dynamics of networks with different connectivity patterns. As illustrated in Figure 1, ACNs demonstrate layer-wise training patterns in which early layers receive significantly stronger gradients during the initial stages of training, in contrast to the more uniform gradient distribution observed in Residual Networks. Next, we empirically demonstrate a broad range of advantages that ACN-learned representations offer compared to residual or feedforward architectures, including: enhanced information compression, superior generalization, reduced catastrophic forgetting, and efficient transferability. Our contributions can be summarized as:
  • We propose a novel architecture, auto-compressing networks (ACNs), that performs auto-compression organically through architectural design, addressing parameter redundancy in deep, overparameterized neural networks—a prevalent issue in modern architectures.
@@
  • We provide a detailed analysis of the gradient dynamics of ACNs, along with residual and feedforward networks, shedding light on their distinct behaviors and arguing that different connectivity patterns result in distinct learned representations.
  • We implement ACNs in fully connected and transformer-based architectures and find experimentally that they achieve similar or superior performance compared to residual baselines, while 30-80% of the top layers become effective identity mappings, as all relevant information is concentrated in the bottom layers. We highlight that this approach is practical with current hardware and does not require specialized software.
@@
  • We argue that auto-compression offers a natural pathway to continual learning by preserving unused parameters for new tasks and utilizing different parameters for different tasks. We empirically validate this by showing that ACNs reduce catastrophic forgetting by up to 18% compared to residual networks in continual learning by preserving capacity for new unseen tasks.
  • We demonstrate that ACNs outperform regularization-based approaches (relying on intermediate losses) at generalization and transfer learning without requiring hyperparameter tuning.
  • We pair ACNs with widely-used baseline pruning techniques, demonstrating that their organically compressed representations significantly amplify the effectiveness of traditional compression methods, achieving superior levels of sparsity compared to residual architectures.
  • We show how long-connection architectures naturally integrate feedback mechanisms, layer-wise training and developmental pruning aspects, potentially offering computational models that could inform our understanding of neurocognitive information processing while advancing more efficient artificial neural networks.
@@ 2 Auto-Compressing Networks
The core idea behind ACNs  [^1] Code for the paper is available here. is to force each layer to produce features that are directly useful for prediction. In this manner, when the last layers are pruned, earlier layers can be used for prediction directly without the need for further fine-tuning. Concretely, we propose replacing the residual short connections with long connections, as described in Eq. 1 and shown in Table 1 for a network of depth $L$ [^2] We note that a classification head can be built on top of $y$. :
@@
In ACNs the output of each layer  [^3] Also the embedded input, represented with $x_{0}$ in equation 1. is directly connected to the output of the network, and thus is directly optimized by the objective function during gradient descent training. Furthermore, the number of possible shortcuts is equal to the number of layers $L$. We find this simplification maintains the improved signal flow that shortcut connections provide, while also introducing the ability to detect potential parameter redundancy in the architecture [^4] A careful reader may observe that long connections are a strict subset of the $2^{L}$ shortcut connections in residual networks.. We note that ACNs differ structurally from other models employing long connections, such as DenseNets (Huang et al., 2017b) and DenseFormer (Pagliardini et al., 2024), which are residual networks variants. These two models connect each layer to all preceding layers whereas ACNs connect each layer only to the output, leading to a distinct structural design [^5] We mention more details about these models in Section 10.
@@ 2.1 Gradient Propagation Across Network Architectures
To understand how different neural network architectures behave during training, we analyze their gradient flow characteristics. In this section, we examine and compare the forward and backward pass (gradient flow) dynamics of three architectures: traditional feedforward networks (FFN), residual networks (ResNet), and the proposed auto-compressing networks (ACN), based on the equations of Table 1. See Appendix B for a detailed derivation of the gradient equations for 1D linear neural networks.
Figure 1: (left) ACNs vs Residual Networks gradient flow across layers during training for ViT architecture (Dosovitskiy, 2020) on ImageNet-1K, showcasing implicit layer-wise training and information concentration on the bottom layers for ACNs. On the other hand, Residual Networks show higher gradient norms (information concentration) in early and deep layers, while middle layers receive significantly lower gradients (suggesting potential parameter redundancy). (right) ACNs vs Residual Networks incremental performance contribution across layers during training for MLP-Mixer architecture (Tolstikhin et al., 2021) on CIFAR-10 Krizhevsky (2009), revealing auto-compression by gradual layer-wise training in ACNs (task-learning starts from shallow layers and gets "pushed" to deeper layers to maximize performance). In Residual Networks, as shown in the Figure task learning happens in the 2-3 final layers.
@@ 2.2 Emergent Gradient Paths
The forward and backward components: As shown in Table 1, each gradient $w_{i}$ (see $\frac{\partial y_{*}}{\partial w_{i}}$ in the 3rd column of the respective table) decomposes into forward and backward terms. The forward term consists of the forward propagated signal up to layer $i$ and determines gradient and forward propagation stability (whether the signal vanishes or explodes), while the backward term influences learning, containing information coming from the loss and traversing subsequent layers. For backward paths, 1D FFNs contain a single path, while 1D ACNs have $L-i+1$: one direct path using the layer’s own long connection to the output, plus $L-i$ additional paths where gradient flows from each subsequent layer’s long connection and back through the network. 1D ResNets have $2^{L-i}$ paths since at each layer there are two options: flow through the network or follow the residual connection. It is also worth observing that ACNs feature a forward term identical to FFNs for intermediate layers (single path), while their backward components is closer to ResNets since it consists of multiple paths. Finally, it is worth mentioning that the FFN path is a subset of the ACN paths, which in turn are contained in the ResNet paths, so, for the set of paths $B$ of the backward term one may write:
$B_{FFN}\subset B_{ACN}\subset B_{ResNet}.$ (2)
Decomposition of the backward term:
The backward component (Full Gradient - FG) can be further decomposed into a Network-mediated Gradient (NG) component that is scaled by network weights and backpropagates information through (a subset of) the network and a Direct Gradient (DG) component that directly connects from the output to each layer, shown as the term "$1$" in the backpropagation equations of ACNs and ResNets in Table 1. This direct path  [^6] This backward path has been previously explored as an alternative to traditional backpropagation and is typically referred to as Direct Feedback Alignment (DFA) in the literature (Nøkland, 2016; Refinetti et al., 2021). acts as an information super highway, especially early in training where weights are typically initialized close to zero, informing each layer directly how to contribute towards lowering the optimization objective. Finally, the DG contribution is more significant for ACNs compared to ResNets, due to ACNs’ linear (rather than exponential) total gradient path count.
@@
Main Claim:
We postulate that: 1) a strong DG component coupled with a weaker feed-forward signal leads implicitly to efficient layer-wise training, and 2) architecturally-induced layer-wise training results inadvertently in a form of structural learning where information is naturally pushed to early layers, i.e., later layers will become redundant (effectively identity mappings) if the earlier layers can already solve for the task. We refer to this new class of networks as auto-compressors since they naturally “shed” their redundant layers during backpropagation simply via architectural design. These claims are experimentally validated in the rest of the paper.
2.3 A Toy Demonstration
Following the theoretical analysis of gradient dynamics, we validate our key claims next through a series of experiments. First, to demonstrate how ACNs naturally compress information into early layers during training, we perform a simple toy experiment involving a $1$D linear feedforward network with three layers and weights $w_{1}$, $w_{2}$ and $w_{3}$ for each layer, respectively. The dataset comprises pairs drawn from the function $y=2x$. We consider two architectures: the first employs residual (short) connections, while the second utilizes long connections (ACN). For this toy problem, both the residual and ACN networks should ideally require only a single layer to successfully accomplish the task, i.e., $w_{1}=1$, $w_{2}=0$, $w_{3}=0$ is the most efficient solution:
[…]
We perform this simple training experiment multiple times ($N=1000$), each time generating $1000$ examples of input $x$ with values between $-10$ and $10$. Both models are trained for $300$ epochs and the weights are initalized with values around zero either uniformly in $[-1,1]$ or following a normal distribution. Fig. 2 illustrates the distribution of learned $w_{1}$ values for both architectures.
Figure 2: Histogram (1000 runs) of the weight of the first layer $w_{1}$ of a $1$D linear feedforward residual network with three layers solving $y=2x$, utilizing either: (a) residual connections or (b) long connections (ACN).
For residual architectures in (a) we observe a pretty wide distribution of $w_{1}$ values centered around $0.26$; indeed $w_{1}=w_{2}=w_{3}\approx 0.26$ is a valid solution of Eq. 3 . Thus, in this example, residual networks have the tendency to utilize all layers equally, even if a sparse solution exists. ACNs, however, typically converge to solutions where $w_{1}$ is close to 1, allowing correct predictions from the first layer [^7] Note that initialization plays a crucial role, as $w_{1}$ sometimes converges near $-1$ for ACNs. Further, the solution that ACN converges at is $w_{1}\approx 0.9$, $w_{2}\approx 0.11$, $w_{3}\approx 0$, so it just gets pretty close to the most efficient solution but it does not achieve perfect compression.. So in this simple example, long connections in ACNs induce implicit depth regularization, guiding the network toward sparse solutions.
The careful reader will observe that this behavior arises from the weight asymmetry in Eq. 4, where there are three terms that include $w_{1}$, two terms with $w_{2}$ and a single term with $w_{3}$. In terms of derivative flow, the DG component for each layer is pretty strong compared to NG. Compare this with the weight symmetric Eq. 3, where the NG component dominates. The residual network naturally converges to the $w_{1}=w_{2}=w_{3}$ symmetric solution [^8] Another way to interpret Eq. 4 is that we have superimposed four feedforward networks: the identity network, a single layer, two layer and three layer network. The important tweak here is that their weights are tied, e.g., $w_{1}$ is common for all network depths, which biases the network towards a shallow solution..

### v2

@@ Abstract
Deep neural networks with short residual connections have demonstrated remarkable success across domains, but increasing depth often introduces computational redundancy without corresponding improvements in representation quality. We introduce Auto-Compressing Networks (ACNs), an architectural variant where additive long feedforward connections from each layer to the output replace traditional short residual connections. By analyzing the distinct dynamics induced by this modification, we reveal a unique property we coin as auto-compression—the ability of a network to organically compress information during training with gradient descent, through architectural design alone. Through auto-compression, information is dynamically "pushed" into early layers during training, enhancing their representational quality and revealing potential redundancy in deeper ones. We theoretically show that this property emerges from layer-wise training patterns present in ACNs, where layers are dynamically utilized during training based on task requirements. We also find that ACNs exhibit enhanced noise robustness compared to residual networks, superior performance in low-data settings, improved transfer learning capabilities, and mitigate catastrophic forgetting suggesting that they learn representations that generalize better despite using fewer parameters. Our results demonstrate up to 18% reduction in catastrophic forgetting and 30-80% architectural compression while maintaining accuracy across vision transformers, MLP-mixers, and BERT architectures. These findings establish ACNs as a practical approach to developing efficient neural architectures that automatically adapt their computational footprint to task complexity, while learning robust representations suitable for noisy real-world tasks and continual learning scenarios.
@@ 1 Introduction
Deep learning has achieved significant breakthroughs across diverse tasks and domains (krizhevsky2012imagenet, ; lecun2015deep, ; brown2020languagemodelsfewshotlearners, ); however, it still lacks the flexibility, robustness, and efficiency of biological networks. Modern models rely on deep architectures with billions of parameters, leading to high computational, storage, and energy costs. Architecturally, these large models are primarily characterized by short residual connections (he2016deep, ), a design initially developed to enable robust training of deep neural networks via backpropagation by providing stable gradient flow through these exceptionally deep networks. These skip connections establish a network topology where multiple information pathways are created (veit2016residual, ), resulting in an ensemble-like behavior that delivers more efficient training and superior generalization compared to traditional feedforward networks.
Historically, since the emergence of Highway Networks (srivastava2015highway, ), which first proposed additive skip connections researchers have explored numerous architectural variations. Residual Networks (ResNets) (he2016deep, ) removed learned gating functions and adopted direct identity skip connections becoming the industry standard. DenseNets (huang2017densely, ) utilized feature concatenation instead of addition, while FractalNets (larsson2017fractalnet, ) introduced a recursive tree-like architecture combining subnetworks of multiple depths to further enrich feature fusion. More recent works include learned weighted averaging across layer outputs (pagliardini2024denseformer, ), application of attention mechanisms across block outputs (elnokrashy2022depth, ) and denser connectivity patterns between network nodes (zhu2025hyperconnections, ). Other works have explored adding scalars to either the residual or block stream to improve performance, training stability and representation learning (savarese2016learning, ; bachlechner2021rezero, ; zhang2024residual, ; fischer2023field, ). In neural machine translation, researchers have drawn inspiration from both vision and language domains to combine information from different layers, enabling richer semantic and spatial propagation throughout the network (dou-etal-2018-exploiting, ; yu2018deep, ).
@@
Table 1: Connectivity (2D case), Forward and Backward Propagation (1D linear case) for FFN, ResNet, and ACN architectures.
@@
While Residual Networks, as discussed, offer efficient and effective training especially in very deep architectures, there are various works that explore behaviors of these architectures that may harm generalization. In (veit2016residual, ; sun2025transformer, ) it is shown that these architectures exhibit a notable resilience to layer dropping and permutation, which could indicate potential redundancy of some layers (i.e. removing a layer does not affect the network). In (huang2016deep, ), it was further observed that dropping subsets of layers during training can reduce overfitting and improve generalization. In a related study (alain2016understanding, ), the authors showed that introducing skip connections between layers can lead to parts of the network being effectively bypassed and under-trained. In (zhang2024residual, ), they show that unscaled residual connections can harm generalization capabilities in generative representation learning. Finally, more recently, research has revealed substantial parameter redundancy and inefficient parameter usage in large-scale foundation models, particularly within their deeper layers (gromov2025the, ; csordas2025language, ). All these observations can be unified under the perspective that, although residual architectures facilitate training via multiple signal pathways, these same pathways can sometimes act as shortcuts that cause certain components to be either underutilized or prone to overfitting—ultimately limiting effective generalization.
@@
Inspired by these observations, in this work we explore whether we can design alternative architectures that preserve the key advantages of Residual Networks—such as multiple signal pathways and efficient gradient flow—while addressing issues like redundancy and shortcut overuse, ultimately enabling more efficient parameter utilization and improved representation learning. To this end, we propose an architectural variant where additive long feedforward connections from each layer to the output replace traditional short residual connections as shown in Table 1, introducing Auto-Compressing Networks (ACNs). ACNs showcase a unique property we coin as auto-compression—the ability of a network to organically compress information during training with gradient descent, through architectural design alone, dynamically pushing information to bottom layers, enhancing their representational quality, and naturally revealing redundant in deeper layers. We theoretically investigate the emergence of this property by analyzing the gradient dynamics of networks with different connectivity patterns. As illustrated in Figure 1, ACNs demonstrate layer-wise training patterns in which early layers receive significantly stronger gradients during the initial stages of training, in contrast to the more uniform gradient distribution observed in Residual Networks. Next, we empirically demonstrate a broad range of advantages that ACN-learned representations offer compared to residual or feedforward architectures, including: enhanced information compression, superior generalization, reduced catastrophic forgetting, and efficient transferability. Our contributions can be summarized as:
  • We introduce Auto-Compressing Networks (ACNs), an architecture that organically compresses information into a subset of the full network’s layers, through architectural design alone.
@@
  • We provide a detailed analysis of the gradient dynamics of ACNs, along with residual and feedforward networks, shedding light on their distinct behaviors and arguing that different connectivity patterns result in unique training regimes that drive distinct learned representations.
  • We implement ACNs in fully connected and transformer-based architectures, finding that they match or outperform residual baselines, while 30–80% of top layers effectively become redundant as information concentrates in the lower layers. Notably, ACNs are hardware-friendly and require no specialized software.
@@
  • ACNs reduce catastrophic forgetting by up to 18% compared to residual networks in continual learning by preserving capacity for new unseen tasks.
  • ACNs outperform regularization-based approaches at transfer learning without requiring hyperparameter tuning.
@@ 2 Auto-Compressing Networks
In ACNs  [^1] Code for the paper is available here., residual short connections are replaced with long feedforward connections, as described in Eq. 1 for a network of depth $L$, and shown  [^2] The connections shown for ResNet (red) are implicit, resulting from the residual summation. A detailed example is provided in Appendix D. in Table 1:
@@
In ACNs the output of each layer is directly connected to the output of the network, and thus is directly optimized by the objective function during gradient descent training. Furthermore, the number of possible shortcuts is equal to the number of layers, $L$. We find this simplification maintains the improved signal flow that shortcut connections provide, while also introducing the ability to detect potential parameter redundancy in the architecture [^3] Note that long connections are a strict subset of the $2^{L}$ shortcut connections in residual networks.. We note that ACNs differ structurally from other models employing long connections, such as DenseNets (huang2017densely, ) and DenseFormer (pagliardini2024denseformer, ), which are residual networks variants, as previously discussed. These models connect each layer to all preceding layers whereas ACNs connect each layer only to the output, leading to a distinct structural design. An overview of these models along with an extended literature review is provided in Appendix A. For completeness, we also provide a direct, fair comparison against other residual variants, such as DenseFormers and DenseNets, in Appendix E.
@@ 2.1 Gradient Propagation Across Network Architectures
In this section, we examine and compare the forward and backward pass (gradient flow) dynamics of three architectures: traditional feedforward networks (FFN), residual networks (ResNet), and the proposed auto-compressing networks (ACN), based on the equations of Table 1. See Appendix C for a detailed derivation of the gradient equations for 1D linear neural networks.
Figure 1: (left) ACNs vs Residual Networks gradient flow across layers during training for MLP-Mixer architecture (tolstikhin2021mlp, ) on CIFAR-10 krizhevsky2009learning , showcasing implicit layer-wise training and information concentration on the bottom layers for ACNs. (right) ACNs vs Residual Networks incremental performance contribution across layers for ViT architecture (dosovitskiy2021an, ) on ImageNet-1K, revealing auto-compression by gradual layer-wise training in ACNs.
@@ 2.2 Emergent Gradient Paths
The forward and backward components: As shown in Table 1, each gradient $w_{i}$ (see $\frac{\partial y_{*}}{\partial w_{i}}$ in the 3rd column of the respective table) decomposes into forward and backward terms. The forward term determines gradient and forward propagation stability (whether the signal vanishes or explodes), while the backward term influences learning. For backward paths, 1D FFNs contain a single path, while 1D ACNs have Li+1L-i+1: one direct path using the layer’s own long connection to the output, plus LiL-i additional paths where gradient flows from each subsequent layer’s long connection and back through the network. 1D ResNets have $2^{L-i}$ paths since at each layer there are two options: flow through the network or follow the residual connection. It is also worth observing that ACNs feature a forward term identical to FFNs for intermediate layers (single path), while their backward components is closer to ResNets since it consists of multiple paths.
The Backward component (Full Gradient - FG) can be further decomposed into a Network-mediated Gradient (NG) component that is scaled by network weights and backpropagates information through (a subset of) the network and a Direct Gradient (DG) component that directly connects from the output to each layer, shown as the term "11" in the backpropagation equations of ACNs and ResNets in Table 1444This backward path has been previously explored as an alternative to traditional backpropagation and is typically referred to as Direct Feedback Alignment (DFA) in the literature (nokland2016direct, ; refinetti2021align, ). This direct path acts as an information super highway, especially early in training where weights are typically initialized close to zero, informing each layer directly how to contribute towards lowering the optimization objective. Finally, the DG contribution is more significant for ACNs compared to ResNets, due to ACNs’ linear (rather than exponential) total gradient path count.
@@
Thus, we postulate that: 1) a strong DG component coupled with a weaker feed-forward signal leads implicitly to efficient layer-wise training, and 2) architecturally-induced layer-wise training results inadvertently in a form of structural learning where information is naturally pushed to early layers, i.e., later layers will become redundant (effectively identity mappings) if the earlier layers can already solve for the task. We refer to this new class of networks as auto-compressors since they naturally “shed” their redundant layers during backpropagation simply via architectural design. These claims are experimentally validated in the rest of the paper.